The activation energy for a given reaction is 50.3 kJ/mol. If the rate constant for the reaction is 0.471 hr-1 at 245 °C , at what temperature, in °C, will the reaction have a rate constant of 0.0107 hr-1 ?
Answer #1

ln (K2/K1) = (Ea/R) [1/T1 - 1/T2]

ln (0.0107 / 0.471) = (50.3 x 1000 / 8.314) [1/518 - 1/T2]

- 3.78 = 50300 / 8.314 [0.0019 - 1/T2]

- 0.00062 = 0.0019 - 1/T2

1/T2 = 0.00252

T2 = 396.8 K

T2 = 123.80C

answer = 124 0C